an object weight 712.0 N is is being pushed along a horizontaL surface by a 440.0 N force. what is the mass of the object? how would you find that? what is newton force? how do you use it in this situation? is it really needed to find the mass of the object?
1
Expert's answer
2012-11-27T10:43:40-0500
F=712 N
F=mg
m=712/9.8=72.6 kg
Newton force -F-mg=712-440=272. We don't need to find mass, because we don't use it